Red Royal Packaging stock is expected to pay a dividend of 3.94 dollars in 1 year and a dividend of 2.8 dollars in 2 years. The stock is expected to be priced at 98.73 dollars in 1 year and at 108.76 dollars in 2 years. What is the current price of Red Royal Packaging stock? The stock’s dividend is paid annually and the next dividend is expected in 1 year.
Answer #1

Expected Dividend in 1 year, D1 = $3.94
Expected Dividend in 2 years, D2 = $2.80
Expected Price in 1 year, P1 = $98.73
Expected Price in 2 years, P2 = $108.76

Let required return be r%

P1 = (D2 + P2) / (1 + r)
$98.73 = ($2.80 + $108.76) / (1 + r)
$98.73 = $111.56 / (1 + r)
1 + r = 1.1300
r = 0.13 or 13%

Current Price, P0 = (D1 + P1) / (1 + r)
Current Price, P0 = ($3.94 + $98.73) / (1 + 0.13)
Current Price, P0 = $102.67 / 1.13
Current Price, P0 = $90.86